# Generative Adversarial Networks for High-Dimensional Item Factor Analysis: A Deep Adversarial Learning Algorithm

### v2

@@ 4 Experiment Results of the Comparison between AVB and VAE
This section describes the results of the empirical study and simulation studies comparing the IWAVB and the IWAE by Urban and Bauer, (2021). IWAE was shown to be more efficient (Urban and Bauer,, 2021) than the popular MH-RM algorithm (Cai,, 2010), while IWAVB achieved similar finite sample performance as IWAE in normal latent variables cases but outperformed it when handling more complex latent variables distribution. All experiments were conducted on a Kaggle computational environment equipped with an Intel(R) Xeon(R) CPU @ 2.00 GHz (4 vCPUs) and two Tesla T4 GPUs (each with 16 GB of memory).
@@ 4.1 Empirical Study: Big-Five Personality Questionnaire
To enhance the interpretability of the loading matrix, we applied the Geomin oblique rotation method (Yates,, 1987) to the loading matrices produced by both algorithms. This method allows factors to correlate, which typically captures the underlying structure more faithfully when the factors are not entirely independent. As shown in Figure LABEL:fig:load_avb_vae, the heat maps of the Geomin-rotated loadings from the reference models of both methods reveal a highly similar and theoretically expected five-factor configuration.
Denote one of the estimates by the reference models is $\xi_{ref}$, and the approximate root-mean-square error (RMSE) across repetitions is defined as follows:
@@ 4.2 Simulation Study: Latent Variables Following Normal Distribution
This study aimed to evaluate the performance of the IWAVB method with respect to parameter recovery and computational efficiency. The evaluation was based on a confirmatory analysis of a graded response model with $P=5$ latent factors, $N=500$ respondents, $M=50$ items, and $C_{j}=5$ response categories for $j=1,\dots,M=50$. Each factor was assumed to load on ten items, and each item corresponded to only one factor. Additionally, all loading values were constrained to be positive. The true correlation matrix of the factors was sampled from the Lewandowski-Kurowicka-Joe (LKJ) distribution (Lewandowski et al.,, 2009), a commonly used prior for positive definite symmetric matrices with unit diagonals in Bayesian hierarchical modeling.
where $\hat{\xi}^{(a)}$ represents the estimate from replication $a$.
After some experimentation, the neural network configuration for IWAVB method was decided as one hidden layer with 128 neurons for the encoder and two hidden layers with 256 and 128 neurons for the discriminator. Meanwhile, the base learning rates for the IWAVB method were set to $\eta_{1}=0.01$ for the discriminator network and $\eta_{2}=0.001$ for the encoder network. The inference model in the IWAE method had one hidden layer with 100 neurons and the base learning rate was $\eta=0.001$.
Figure 8 shows box plots of the MSEs for item parameters (loadings and intercepts), and factor correlations for the IWAVB and IWAE methods with 15, 25, or 35 importance-weighted (IW) samples. Both methods achieved low MSEs for estimating item loadings and factor correlations. While the MSE for item intercepts was relatively higher, it still averaged below 0.03, which is satisfactory. Although the IWAE method slightly outperformed IWAVB in estimating item intercepts and factor correlations, the differences were minimal. Notably, IWAVB achieved a smaller MSE for item loadings. Both methods showed marginal improvements in estimation accuracy as the number of IW samples increased.
Figure 8: Parameter MSE comparison for IWAVB and IWAE methods based on 100 replications of simulation with latent variables following normal distribution.
This simulation study aimed to evaluate a more general case where the latent variables follow a multimodal distribution. A graded response model was configured with $P=1$ latent factor, $N=500$ respondents, $M=20$ items, and $C_{j}=3$ response categories for $j=1,\dots,M=20$. The true latent variables were sampled from a mixture of three Gaussian distributions: $\mathcal{N}(-1.5,0.5)$, $\mathcal{N}(0,0.5)$, and $\mathcal{N}(1.5,0.5)$, with respective weights of $0.4$, $0.2$, and $0.4$, as shown in Figure 10. This resulted in a distribution with multiple modes, in contrast to the unimodal structure of a normal distribution. For the item parameters, 20 true loading values were generated from a log-normal distribution with a mean of zero and variance of $0.5$. True intercepts were sampled using a similar procedure as described in the previous simulation study. Additionally, a three-dimensional case was examined by configuring a GRM with $P=3$ latent factor, $N=5000$ respondents, $M=45$ items, and $C_{j}=3$ per item. Each item also corresponded to only one factor. In this instance, the true latent variables were generated from a mixture of three trivariate Gaussian distributions with mean vectors of the form $(\mu,\mu,\mu)$, where $\mu$ takes on the values $-1.5$, 00, and $1.5$, and with a fixed correlation matrix sampled from the LKJ distribution. The true loading and intercept values were generated using analogous procedures as described previously.
Figure 10: Multimodal Distribution of Latent Variables.
We used the following base learning rates for the discriminator and encoder in the IWAVB method: $(\eta_{1}=0.003,\eta_{2}=0.0002)$ for $P=1$ and $(\eta_{1}=0.002,\eta_{2}=0.0005)$ for $P=3$, while for the inference model (one hidden layer with 100 neurons) in the IWAE method, both cases utilized $\eta=0.0005$. The configuration of neural networks was same as the previous study. The number of IW samples was set to $R=30$ for both methods. Using these configurations, Figure 11 presents box plots comparing the MSEs for item loadings and intercepts estimated by the IWAVB and IWAE methods. The IWAVB method consistently achieved lower MSEs across both parameter estimates compared to IWAE.
Figure 12 illustrates the box plots for computation time and approximate log-likelihood in this simulation study. Although the IWAVB method required more computational time, it consistently achieved a higher approximate log-likelihood compared to the IWAE method.
Figure 12: Approximate log-likelihood and computation time comparison for IWAVB and IWAE methods across 100 replications of simulation with latent variables following multimodal distribution.
@@ 5 Discussion
We conducted three comparative studies between IWAE and IWAVB to evaluate their performance. The empirical study, based on the large-scale Big-Five personality factors dataset, yielded parameter estimate results comparable to the IWAE method while achieving higher model likelihood, indicating a better fit to the data. In the first simulation study using a normal distribution for latent variables, IWAVB achieved similarly low mean squared error and higher log-likelihood compared to IWAE, demonstrating its efficacy in standard conditions. Furthermore, in a more general and realistic scenario where the latent variables follow a multimodal distribution, IWAVB outperformed IWAE in the estimation of all parameters and log-likelihood, highlighting its robustness and superior performance in complex and real-world settings.
@@
Nonetheless, the proposed IWAVB method is not without limitations that need further improvement. First, exploring different neural network specification—such as altering the number or size of hidden layers, or integrating more advanced architectures such as attention—may prove beneficial for diverse tasks, yet there is no standardized IFA extension nor well-defined criterion to guide these extensions. Second, as Ma et al., (2024) demonstrate for IWAE, it is reasonable to suspect the extension version, IWAVB, might also struggle to achieve precise item parameter estimates when the respondent sample size is relatively small, indicating an ongoing opportunity to further improve small sample performance for VI-based methods.
Despite these limitations, the concept of leveraging neural networks to represent complex probabilistic models remains promising in psychometrics and beyond. In particular, because the decoder in the proposed AVB method currently relies on a simple sigmoid-activated linear function, more sophisticated neural architectures could improve parameter estimation and provide methodological foundations to extend IFA to accommodate multimodal data inputs, such as modeling image and texts (Cheng et al.,, 2019). Meanwhile, AVB provides a framework that bridges GANs with Marginal Maximum Likelihood, thereby enabling the adaptation of numerous GAN variants, such as Deep Convolutional GANs (Radford,, 2015), Wasserstein GANs (Arjovsky et al.,, 2017), and CycleGANs (Zhu et al., 2017a, ), to potentially improve estimation in IFA and beyond.

### v3

@@ 4 Experiment Results of the Comparison between AVB and VAE
This section describes the results of the empirical study and simulation studies comparing the IWAVB and the IWAE by Urban and Bauer, (2021). IWAE was shown to be more efficient (Urban and Bauer,, 2021) than the popular MH-RM algorithm (Cai,, 2010), while IWAVB achieved similar finite sample performance as IWAE in normal latent variables cases but outperformed it when handling more complex latent variables distribution. The algorithms were implemented in PyTorch 2.6.0. All experiments were conducted on a Kaggle computational environment equipped with an Intel(R) Xeon(R) CPU @ 2.00 GHz (4 vCPUs) and two Tesla T4 GPUs (each with 16 GB of memory).
4.1.1 Comparison with IWAE: Importance-Weighting Technique
This simulation study compared the IWAVB and IWAE methods in terms of parameter recovery and computational efficiency. Because both methods employ importance weighting, comparisons were conducted using various numbers of IW samples. The evaluation was based on a confirmatory analysis of a graded response model with $P=5$ latent factors, $N=500$ respondents, $M=50$ items, and $C_{j}=5$ response categories for $j=1,\dots,M=50$. Each factor was assumed to load on ten items, and each item corresponded to only one factor. Additionally, all loading values were constrained to be positive. The true correlation matrix of the factors was sampled from the Lewandowski-Kurowicka-Joe (LKJ) distribution (Lewandowski et al.,, 2009), a commonly used prior for positive definite symmetric matrices with unit diagonals in Bayesian hierarchical modeling.
where $\hat{\xi}^{(a)}$ represents the estimate from replication $a$. Bias is also computed by removing the square for each summand in the Equation (21).
After some experimentation for hyper-parameter tunning, the neural network configuration for IWAVB method was decided as one hidden layer with 128 neurons for the encoder and two hidden layers with 256 and 128 neurons for the discriminator. Meanwhile, the base learning rates for the IWAVB method were set to $\eta_{1}=0.01$ for the discriminator network and $\eta_{2}=0.001$ for the encoder network. The inference model in the IWAE method had one hidden layer with 100 neurons and the base learning rate was $\eta=0.001$.
Figure 5 shows box plots of the MSEs and bias for item parameters (loadings and intercepts), and factor correlations for the IWAVB and IWAE methods with 15, 25, or 35 importance-weighted (IW) samples. Both methods achieved low MSEs for estimating item loadings and factor correlations. While the MSE for item intercepts was relatively high, it still averaged below 0.03, which is satisfactory. Although the IWAE method slightly outperformed IWAVB in estimating item intercepts and factor correlations, the differences were minimal. Notably, IWAVB achieved a smaller MSE for loadings. Bias values for both methods were consistently low, with loading estimates showing the lowest bias among the three parameter types. Both methods showed marginal improvements in estimation accuracy as the number of IW samples increased.
Figure 5: Parameter MSE and bias comparison for IWAVB and IWAE methods based on 100 replications of simulation with 5-dimensional latent variables following normal distribution.
4.1.2 Comparison with MH-RM: Higher Dimensionality and Larger Sample Size
To assess performance on more demanding problems, this study extended the comparison to include the established MH-RM algorithm. We used the MH-RM implementation from the mirt (Version 1.44.0; Chalmers,, 2012). Since its core functions are written in C++, it provides a relevant benchmark against the C++ backend of PyTorch. The comparison was based on two large-scale confirmatory analyses with the GRM, both using a sample size of $N=5000$ respondents: (1) a 7-dimensional model with $P=7$ latent factors and $M=70$ items; (2) a 10-dimensional model with $P=10$ latent factors and $M=100$ items. In both scenarios, each item had $C_{j}=5$ response categories. The factor structure remained simple, with each factor loading on ten unique items. The process for generating the true latent variables, item parameters, and the factor correlation matrix (from the LKJ distribution) was similar as the previous simulation.
Hyperparameter tuning indicated that the neural network configuration from the previous simulation remained effective for the 7-dimensional IWAVB model. For the 10-dimensional IWAVB model, one hidden layer with 170 neurons for the encoder and two hidden layers (400 and 200 neurons, respectively) for the discriminator provided optimal results. For IWAE, a single hidden layer with 100 neurons remained sufficient. Learning rates were unchanged from the prior simulations. The number of IW samples was set to 50 for IWAVB and IWAE methods to ensure stable performance.
Figure 7 presents box plots comparing MSEs for loadings, intercepts, and factor correlations across IWAVB, IWAE, and MH-RM. All three methods demonstrated low MSE values, with negligible differences between methods.
Figure 7: Parameter MSE comparison for IWAVB, IWAE and MHRM methods based on 100 replications of simulation with 7-dimensional and 10-dimensional latent variables following normal distribution.
Table 1 compares computation times among the three methods. Due to the lack of GPU support in mirt, MH-RM required substantially more computation time, especially in higher-dimensional cases with larger sample sizes. IWAVB, while still slower than IWAE, significantly outperformed MH-RM in computational efficiency.
Table 1: Computation time comparison for IWAE, IWAVB and MH-RM methods across 100 replications of simulation with 7-dimensional and 10-dimensional latent variables following normal distribution.
IWAE (7-dim) IWAVB (7-dim) MH-RM (7-dim)
Mean computation time (seconds) 76.66 (SD=12.07) 159.99 (SD=26.63) 662.93 (SD=25.48)
IWAE (10-dim) IWAVB (10-dim) MH-RM (10-dim)
Mean computation time (seconds) 80.64 (SD=14.00) 179.03 (SD=29.85) 1,666.54 (SD=43.71)
Note: SD = Standard Deviation.
This simulation study aimed to evaluate a more general case where the latent variables follow a multimodal distribution. A graded response model was configured with $P=1$ latent factor, $N=500$ respondents, $M=20$ items, and $C_{j}=3$ response categories for $j=1,\dots,M=20$. The true latent variables were sampled from a mixture of three Gaussian distributions: $\mathcal{N}(-1.5,0.5)$, $\mathcal{N}(0,0.5)$, and $\mathcal{N}(1.5,0.5)$, with respective weights of $0.4$, $0.2$, and $0.4$, as shown in Figure 8. This resulted in a distribution with multiple modes, in contrast to the unimodal structure of a normal distribution. For the item parameters, 20 true loading values were generated from a log-normal distribution with a mean of zero and variance of $0.5$. True intercepts were sampled using a similar procedure as described in the previous simulation study. Additionally, a three-dimensional case was examined by configuring a GRM with $P=3$ latent factor, $N=5000$ respondents, $M=45$ items, and $C_{j}=3$ per item. Each item also corresponded to only one factor. In this instance, the true latent variables were generated from a mixture of three trivariate Gaussian distributions with mean vectors taking on the values 1.5-1.5, 0, and $1.5$, and with a fixed correlation matrix sampled from the LKJ distribution. The true loading and intercept values were generated using analogous procedures as described previously.
Figure 8: Multimodal Distribution of Latent Variables.
We used the following base learning rates for the discriminator and encoder in the IWAVB method: $(\eta_{1}=0.003,\eta_{2}=0.0002)$ for $P=1$ and $(\eta_{1}=0.002,\eta_{2}=0.0005)$ for $P=3$, while for the inference model (one hidden layer with 100 neurons) in the IWAE method, both cases utilized $\eta=0.0005$. The configuration of neural networks was same as the first study. The number of IW samples was set to $R=30$ for both methods. Using these configurations, Figure 9 presents box plots comparing the MSEs for item loadings and intercepts estimated by the IWAVB and IWAE methods. The IWAVB method consistently achieved lower MSEs across both parameter estimates compared to IWAE.
Table 2 illustrates the computation time and approximate log-likelihood comparison for two methods in this simulation study. Although the IWAVB method required more computational time, it consistently achieved a higher approximate log-likelihood compared to the IWAE method.
Table 2: Computation time and approximate log-likelihood comparison for IWAE and IWAVB methods across 100 replications of simulation with latent variables following multimodal distribution.
IWAE (1-dim) IWAVB (1-dim)
Mean computation time (seconds) 162.01 (SD=25.19) 181.03 (SD=22.09)
Mean log-likelihood -5,233.35 (SD=80.38) -5,129.92 (SD=102.19)
IWAE (3-dim) IWAVB (3-dim)
Mean computation time (seconds) 217.42 (SD=42.16) 255.93 (SD=45.99)
Mean log-likelihood -125,398.75 (SD=325.64) -111,723.19 (SD=537.28)
Note: SD = Standard Deviation.
@@ 4.3 Empirical Study: Big-Five Personality Questionnaire
To enhance the interpretability of the loading matrix, we applied the Geomin oblique rotation method (Yates,, 1987) to the loading matrices produced by both algorithms. This method allows factors to correlate, which typically captures the underlying structure more faithfully when the factors are not entirely independent. As shown in Figure 12(a), the heat maps of the Geomin-rotated loadings from the reference models of both methods reveal a highly similar and theoretically expected five-factor configuration.
(a) Heat map of factor loadings estimated by IWAE and IWAVB for IPIP-FFM items.
Let $\xi_{ref}$ denote one of the parameter estimates obtained from the reference model, and the approximate root-mean-square error (RMSE) across repetitions is defined as follows:
@@ 5 Discussion
We conducted three comparative studies between IWAE and IWAVB to evaluate their performance. In the first simulation study using a normal distribution for latent variables, IWAVB achieved similarly low mean squared error and higher log-likelihood compared to IWAE, while demonstrating its higher efficiency in standard conditions than established MH-RM algorithm. Furthermore, in a more general and realistic scenario where the latent variables follow a multimodal distribution, IWAVB outperformed IWAE in the estimation of all parameters and log-likelihood, highlighting its robustness and superior performance in complex and real-world settings. The empirical study, based on the large-scale Big-Five personality factors dataset, yielded parameter estimate results comparable to the IWAE method while achieving higher model likelihood, indicating a better fit to the data.
@@
Nonetheless, the proposed IWAVB method is not without limitations that need further improvement. First, as with many GAN-based methods, training can be unstable. This instability arises from the adversarial process, in which the generator and discriminator compete rather than optimize a single objective. Consequently, issues such as mode collapse may occur, where an overly strong discriminator in early training restricts the generator’s ability to explore the latent space (Mi et al.,, 2018). Although adjusting hyperparameters or network architecture can mitigate these problems, our experiments show that IWAVB is highly sensitive to such changes, making the tuning process time-consuming. Furthermore, exploring different neural network specification, such as altering the number or size of hidden layers, or integrating more advanced architectures such as attention, may prove beneficial for diverse tasks, yet there is no standardized IFA extension nor well-defined criterion to guide these extensions.
Despite these limitations, the concept of leveraging neural networks to represent complex probabilistic models remains promising in psychometrics and beyond. In particular, because the decoder currently relies on a simple sigmoid-activated linear function, more sophisticated neural architectures could improve parameter estimation and provide methodological foundations to extend IFA to accommodate multimodal data inputs, such as modeling image and texts (Cheng et al.,, 2019). Meanwhile, AVB provides a framework that bridges GANs with Marginal Maximum Likelihood, thereby enabling the adaptation of numerous GAN variants, such as Deep Convolutional GANs (Radford,, 2015), Wasserstein GANs (Arjovsky et al.,, 2017), and CycleGANs (Zhu et al., 2017a, ), to potentially improve estimation in IFA and beyond.
@@
Appendix
Considering $P_{l}\times P_{l-1}$ matrix $\bm{W}_{0}^{(l)}=[w_{p_{l},p_{l-1},0}]$ and $P_{l}\times 1$ vector $\bm{b}_{0}^{(l)}=(b_{p_{l},0})$, Kaiming initialization is defined as follows:
where $\mathcal{U}(a,b)$ represents the uniform distribution with lower bound $a$ and upper bound $b$.
For item $j=1,\dots,M$, the $P\times 1$ vector of loadings $\bm{\beta}_{j,0}=(\beta_{p,j,0})$ and the $(C_{j}-1)\times 1$ vector of intercepts $\bm{\alpha}_{j,0}=(\alpha_{k,j,0})$ are initialized by: for $p=1,\dots,P$ and $k=1,\dots,C_{j}-1$,
AdamW separates the weight decay term from the gradient updates, which enables better generalization in deep learning models because it explicitly penalizes large weights through a regularization term without conflating it with the optimization step. The parameter update rule in AdamW includes momentum and second-order moment estimation, and is expressed as follows:
$\text{Momentum update: }\bm{m}_{t}=\beta_{1}\bm{m}_{t-1}+(1-\beta_{1})\nabla_{\bm{\omega}}\mathcal{L}(\bm{\omega}),$ (26)
$\text{Second moment update: }\bm{v}_{t}=\beta_{2}\bm{v}_{t-1}+(1-\beta_{2})(\nabla_{\bm{\omega}}\mathcal{L}(\bm{\omega}))^{2},$ (27)
$\text{Weight decay adjustment: }\bm{\omega}_{t+1}=\bm{\omega}_{t}-\eta\frac{\bm{m}_{t}}{\sqrt{\bm{v}_{t}}+\epsilon}-\eta\lambda\bm{\omega}_{t},$ (28)
where $\bm{m}_{t}$ and $\bm{v}_{t}$ are the first and second moment estimates, $\beta_{1}$ and $\beta_{2}$ are exponential decay rates for the first and second moment estimates, $\lambda$ is the weight decay coefficient, and $\epsilon$ is a small constant for numerical stability. The hyperparameters are set to the default values ($\beta_{1}=0.9,\beta_{2}=0.999,\lambda=0.01,\epsilon=1\times 10^{-8}$) in this study, as they are practically recommended (Loshchilov,, 2017; Reddi et al.,, 2019).